# Bottom-up Policy Optimization: Your Language Model Policy Secretly Contains Internal Policies

### v1

@@ A.1 Implementation of BuPO
0: Policy $\pi_{\theta}$, Dataset $\mathcal{Q}$, Target Layer $l$, Internal Steps $s_{\text{inter}}$, Max Steps $S_{\text{max}}$
@@ A.3 The Template for Experiments
We adopt the following template for all experiments involving Qwen models, building upon the Qwen-Math template used for Qwen2.5 (yang2024qwen2) and the Qwen-Nothinking template for Qwen3.
@@
For training the Llama-OctoThinker models, we adopt the original prompt in wang2025octothinker to ensure performance.
@@ A.4 Implementation of Internal Policy Entropy Analysis
In this section, we detail the implementation of internal policy entropy analysis. Our primary objective is to extract internal hidden states during the forward pass. In the main experiments, we evaluate model-generated responses on the MATH test set (hendrycks2021measuring). Entropy is computed at the token level for each layer and module, and then averaged over all generated tokens. We find that the intrinsic reasoning patterns remain stable across different tasks, e.g., commonsense question answering (rein2024gpqa).
@@
$\Delta H^{l}_{\text{Layer}}=\mathcal{H}(\mathbf{H}^{l}_{\text{Layer}})-\mathcal{H}(\mathbf{H}^{l-1}_{\text{Layer}}).$ (12)
@@
$\Delta H^{l}_{\text{ATTN}}=\mathcal{H}(\mathbf{A}^{l})-\mathcal{H}(\mathbf{X}^{l}_{\text{attn}}),$ (13)
@@
$\Delta H^{l}_{\text{FFN}}=\mathcal{H}(\mathbf{F}^{l})-\mathcal{H}(\mathbf{X}^{l}_{\text{ffn}}).$ (14)
@@
Discussion 1: Comparison to the logit lens.
Notably, our definition of internal policy differs from the logit lens approach (nostalgebraist2020logitlens), particularly in how layer normalization (LN) is handled. To clarify this distinction, we provide a systematic comparison between the two formulations in Table 4.
@@ A.5 Implementation of Internal Policy Optimization
In internal policy optimization, namely InterGRPO, at each optimization step, we select a specific internal layer $l$ and optimize its internal policy $\pi_{\text{Layer}}^{l}$, defined as $\pi_{\text{Layer}}^{l}=\text{softmax}(\texttt{LN}(\mathbf{H}^{l})\mathbf{E}_{\text{u}}^{\text{T}})$, where $\mathbf{H}^{l}$ is the hidden states at layer $l$. The gradient flow for internal policy optimization is determined by the residual structure of the Transformer, where the hidden state $\mathbf{H}^{l}$ is a function of all parameters from layer $1$ to $l$, but is independent of parameters in higher layers.
@@
A.6 Implementation of RL
In Sec. 4 and Sec. 5, we conduct reinforcement learning (RL) experiments to evaluate the performance of the proposed algorithm. In this section, we describe the RL training setup in detail. We implement GRPO and other baseline algorithms using the veRL framework (sheng2025hybridflow). Across all algorithms and model variants, we adopt a unified set of hyperparameters, as reported in Table 5, and do not employ entropy regularization or KL-based losses. For PPO, the critic network is trained separately with a learning rate of $1\times 10^{-5}$.
@@ A.6 Implementation of RL
Max response length 7168 tokens (Qwen3) / 3072 (Others)
@@ B.1 Internal Policy Entropy Dynamics for More Models
In this section, we present additional preliminary analyses of internal policy entropy dynamics across a broader set of models. Specifically, we examine different variants of the same backbone, including Base, Instruct, and Mix versions, as well as models trained with supervised fine-tuning (SFT) and reinforcement learning (RL). In addition, we include the DeepSeek-Math model (he2025deepmath) to further enrich the comparative analysis.
@@
We further analyze models that undergo additional training beyond standard pre-training. Specifically, we include DeepSeek-R1-Distill-Qwen-7B, which is further trained from Qwen2.5-Math-7B using distilled responses from guo2025deepseek with SFT, as well as Llama-OctoThinker-3B-Base and Llama-OctoThinker-8B-Base, which are obtained via continued pre-training (i.e., mid-training) based on Llama-3.2-3B-Base and Llama-3.2-8B-Base, respectively (wang2025octothinker). Moreover, Qwen3-4B and Qwen3-4B-Base also show consistent pattern after post-training with RL. Notably, these additional training procedures exhibit only marginal influence on the internal reasoning patterns.
@@
We further analyze the internal reasoning mechanisms of DeepSeek-Math-7B-Base to provide a more comprehensive comparison. As illustrated in Figure 23(a), DeepSeek-Math-7B-Base exhibits a markedly different entropy dynamics: the overall internal policy entropy decreases substantially and converges primarily in the middle layers. This phenomenon is primarily driven by the consistently negative entropy change in the FFN module, i.e., $\Delta H_{\text{FFN}}^{l}<0$, as illustrated in Figure 23(b), particularly in the middle layers.
@@
(a) Continuous Entropy Dynamics of Internal Policy for additional models. The information flows from $\mathbf{H}^{l-1}$ into $\mathbf{A}^{l}$, $\mathbf{F}^{l}$, and finally to the next layer $\mathbf{H}^{l}$.
(b) Entropy Change Dynamics of Internal Policy with more models.
B.2 Pass@$K$ Performance across Datasets

### v2

@@ A.1 Implementation of BuPO
0: Policy $\pi_{\theta}$, Dataset $\mathcal{Q}$, Target Layer $l$, Internal Training Steps $s_{\text{inter}}$, Total Training Steps $S_{\text{max}}$
@@ A.3 The Template for Experiments
We adopt the following template for all experiments involving Qwen models, building upon the Qwen-Math template used for Qwen2.5 (Yang et al., 2024) and the Qwen-Nothinking template for Qwen3.
@@
For training the Llama-OctoThinker models, we adopt the original prompt in Wang et al. (2025b) to ensure performance.
@@ A.4 Implementation of Internal Policy Entropy Analysis
In this section, we detail the implementation of internal policy entropy analysis. Our primary objective is to extract internal hidden states during the forward pass. In the main experiments, we evaluate model-generated responses on the MATH test set (Hendrycks et al., 2021). Entropy is computed at the token level for each layer and module, and then averaged over all generated tokens. We find that the intrinsic reasoning patterns remain stable across different tasks, e.g., commonsense question answering (Rein et al., 2024).
@@
$\Delta H^{l}_{\text{Layer}}=\mathcal{H}(\mathbf{H}^{l})-\mathcal{H}(\mathbf{H}^{l-1}).$ (12)
@@
$\Delta H^{l}_{\text{ATTN}}=\mathcal{H}(\mathbf{A}^{l})-\mathcal{H}(\texttt{LN}(\mathbf{H}^{(2l-2)})),$ (13)
@@
$\Delta H^{l}_{\text{FFN}}=\mathcal{H}(\mathbf{F}^{l})-\mathcal{H}(\texttt{LN}(\mathbf{H}^{(2l-1)})).$ (14)
@@
Discussion: Comparison to the Logit Lens.
Notably, our definition of internal policy differs from the logit lens approach (nostalgebraist, 2020), particularly in how layer normalization (LN) is handled. To clarify this distinction, we provide a systematic comparison between the two formulations in Table 4.
@@ A.5 Implementation of Internal Policy Optimization
In internal policy optimization, namely InterGRPO, at each optimization step, we select a specific internal layer $l$ and optimize its internal policy $\pi_{\text{Layer}}^{l}$, defined as $\pi_{\text{Layer}}^{l}=\text{softmax}(\mathbf{H}^{l}\mathbf{E}_{\text{u}}^{\text{T}})$, where $\mathbf{H}^{l}$ is the hidden states at layer $l$. The gradient flow for internal policy optimization is determined by the residual structure of the Transformer, where the hidden state $\mathbf{H}^{l}$ is a function of all parameters from layer $1$ to $l$, but is independent of parameters in higher layers.
@@
Training Setup.
Based on the findings in Sec. 3, we select Qwen3-4B (Yang et al., 2025) in non-thinking mode for investigation. For the training set, we randomly sample 5k entries from DeepMath-103k (He et al., 2025). We train the models using the verl framework (Sheng et al., 2025). The prompt batch size is 128, with 8 rollouts generated per prompt. The sampling temperature during training is set to 1.0, and the maximum context length is set to 9,216 tokens. We update the model with a mini-batch size of 32 for 300 steps and a learning rate of 1e-6. For Qwen3-4B, the region boundaries identified in Sec. 3.2 lie at layers 6 and 26. Additionally, we focus on aligning the internal policy of the penultimate layer, which is critical for the final prediction. Accordingly, we compare the internal policies $\pi_{\text{Layer}}^{6}$, $\pi_{\text{Layer}}^{26}$, and $\pi_{\text{Layer}}^{35}$ optimized via InterGRPO, against the overall policy $\pi_{\theta}$ optimized with GRPO (He et al., 2025).
A.6 Implementation of Main Experiments
Detailed Training Setup. Specifically, we focus on the Qwen3 series, which maintain a stable EIC pattern, using Qwen3-4B and Qwen3-8B (Yang et al., 2025). Meanwhile, we select Llama-OctoThinker-3B-Base and Llama-OctoThinker-8B-Base from the Llama series, as these models demonstrate improved training behavior after mid-training based on Llama-3.2-Base (Wang et al., 2025b). We implement GRPO and other baseline algorithms using the veRL framework (Sheng et al., 2025). Across all algorithms and model variants, we adopt a unified set of hyperparameters, as reported in Table 5, and do not employ entropy regularization or KL-based losses. For PPO, the critic network is trained separately with a learning rate of $1\times 10^{-5}$.
[…]
@@ A.6 Implementation of Main Experiments
Max response length 7168 tokens (Qwen) / 3072 (Llama)
@@ B.1 Internal Policy Entropy Dynamics for More Models
In this section, we present additional preliminary analyses of internal policy entropy dynamics across a broader set of models. Specifically, we examine different variants of the same backbone, including Base, Instruct, and Mix versions, as well as models trained with supervised fine-tuning (SFT) and reinforcement learning (RL). In addition, we include the DeepSeek-Math model (He et al., 2025) to further enrich the comparative analysis.
@@
We further analyze models that undergo additional training beyond standard pre-training. Specifically, we include DeepSeek-R1-Distill-Qwen-7B, which is further trained from Qwen2.5-Math-7B using distilled responses from Guo et al. (2025) with SFT, as well as Llama-OctoThinker-3B-Base and Llama-OctoThinker-8B-Base, which are obtained via continued pre-training (i.e., mid-training) based on Llama-3.2-3B-Base and Llama-3.2-8B-Base, respectively (Wang et al., 2025b). Moreover, Qwen3-4B and Qwen3-4B-Base also show consistent pattern after post-training with RL. Notably, these additional training procedures exhibit only marginal influence on the internal reasoning patterns.
@@
(a) Residual cosine similarity across different Qwen models
@@
We further analyze the internal reasoning mechanisms of DeepSeek-Math-7B-Base to provide a more comprehensive comparison. As illustrated in Figure 18(b), DeepSeek-Math-7B-Base exhibits a markedly different entropy dynamics: the overall internal policy entropy decreases substantially and converges primarily in the middle layers. This phenomenon is primarily driven by the consistently negative entropy change in the FFN module, i.e., $\Delta H_{\text{FFN}}^{l}<0$, as illustrated in Figure 18(c), particularly in the middle layers.
@@
B.2 How do Internal modules influence the residual stream?
To further understand how internal modules shape the residual stream in Qwen models with progressive reasoning pattern, we analyze residual cosine similarity, which quantifies how each module writes to the residual pathway (Hu et al., 2025b). For a given layer $l$, we compute $\text{cossim}(\mathbf{A}^{l},\mathbf{H}^{l-1})$ for self-attention and $\text{cossim}(\mathbf{F}^{l},\mathbf{H}^{l-1}+\mathbf{A}^{l})$ for the FFN. A cosine similarity near zero indicates writing new, orthogonal features; negative values indicate feature suppression; and positive values indicate amplification of existing features.
The FFN modulates the residual stream in a stage-dependent manner across all models: in lower layers, it injects largely orthogonal features to support exploration; in middle layers, it suppresses vague signals while integrating parametric knowledge in FFN, corresponding to the Integration stage; and in upper layers, it amplifies andintegrates features to drive convergence. Across all models, the final layer exhibits a sharp directional shift, underscoring its critical role in final prediction (Gupta et al., 2025; Agarwal et al., 2025).
(b) Continuous entropy dynamics of internal policy for additional models. The information flows from $\mathbf{H}^{l-1}$ into $\mathbf{A}^{l}$, $\mathbf{F}^{l}$, then to the next layer $\mathbf{H}^{l}$.
(c) Entropy change dynamics of internal policy with more models.
B.3 Pass@$K$ Performance across Datasets